Suppose that 3 out of 4 dentists recommend SparkleSmile toothpaste.
(a) If 10 dentists are randomly selected, find the probability that less than eight would recommend SparkleSmile.

Answer by stanbon(75887)

Suppose that 3 out of 4 dentists recommend SparkleSmile toothpaste.
(a) If 10 dentists are randomly selected, find the probability that less than eight would recommend SparkleSmile.
----------
Binomial problem with p = 3/4 and n = 10
----
P(0<= x <=7) = binomcdf(10,3/4,7) = 0.4744
---------------
Cheers,
Stan H.
-------------